A) When output is 500, a firm's fixed costs are $10,000 and its variable costs are $15,000. The firm's total costs are therefore: $ 5,000. $10,000. $15,000. $25,000.

B) If fixed costs are $960, variable costs are $1,440, and output is 12, then average total cost equals:

$ 80.

$100.

$120.

$200.
Answer #1

A) Firm's fixed cost = 10000

Its variable cost = 15000

Total cost = fixed cost + variable cost = 10000+15000 = 25000

B) Fixed Cost = 960

Variable cost = 1440

output = 12

Total cost = fixed cost + variable cost = 960+1440 = 2400

Average total cost = total cost / output = 2400 / 12 = 200